Solve the quadratic equation using any appropriate method. 4(x+1)^2=64
Answer by 303795(602)

4(x+1)^2=64
Find the square root of each side.
2(x+1)=8 or 2(x+1)=-8
2x+2=8 or 2x+2=-8
2x=6 or 2x=-10
x=3 or x=-5